OpenStudy (anonymous):

According to a survey published by ComPsych Corporation, 54% of all workers read e-mail while they are talking on the phone. Suppose that 20% of those who read e-mail while they are talking on the phone write personal "to-do" lists during meetings. Assuming that these figures are true for all workers, if a worker is randomly selected, determine the following probabilities: a. The worker reads e-mail while talking on the phone and writes personal "to-do" lists during meetings. b. The worker does not write personal "to-do" lists given that he reads e-mail while talking on the phone.

OpenStudy (anonymous):

OK, let the total no. of workers be P okay?

OpenStudy (anonymous):

okay? work with me here.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so The worker reads e-mail while talking on the phone AND writes personal "to-do" lists during meetings the no. of workers who read e-mail while talking on the phone=0.54(P)

OpenStudy (anonymous):

agreed?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now let this no. be Q. Finally, the no. of workers who read e-mail while talking on the phone AND writes personal "to-do" lists during meetings=0.2(Q), right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so, this no. is 0.2(q)=0.2*0.54*P =0.108 P =10.8% of P

OpenStudy (anonymous):

u mean 10.8% of p is the answer for a

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

Now for b.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Note that if 20% of those who read e-mail while they are talking on the phone write personal "to-do" lists during meetings, 80% of them DO NOT write these to do lists. So the required no. is 80% of Q i.e. 80% of 0.54P =0.8*0.54*P =0.432P =43.2% of P.

OpenStudy (anonymous):

done?

OpenStudy (anonymous):

ya..

OpenStudy (anonymous):

cool.

OpenStudy (anonymous):

i m confused

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

what is that 43.2% of p.. u did already put a value of p i.e. 0.54

OpenStudy (anonymous):

and plz do c too

OpenStudy (anonymous):

No, it's 80% of Q Q is nothing but 0.54P.

OpenStudy (anonymous):

i got one last question for u... if u don't mind...

OpenStudy (anonymous):

sure, but did you understand b?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

ok, if the line containing "i.e" is line 1, which line do you find confusing?

OpenStudy (anonymous):

instead of writing 0.432p can i simply write 0.432

OpenStudy (anonymous):

i guess so, if it confuses you.

OpenStudy (anonymous):

will that be right

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

and c

OpenStudy (anonymous):

post it.

OpenStudy (anonymous):

Jessica Salas, president of Salas Products, is reviewing the warranty policy for her company's new model of automobile batteries. Life tests performed on a sample of 100 batteries indicated that an average life of 75 months, a standard deviation of 5 months, and a bell shaped battery life distribution. (a) between what two numbers would 95% of the batteries will fall? (b) between what two numbers would 99.7% of the batteries will fall?

OpenStudy (anonymous):

this isn't c :P Post it as a new question.

OpenStudy (anonymous):

c) The worker doesnot write personal "to-do" lists and does read e-mailwhile talking on phone.

OpenStudy (anonymous):

Hey dude!!

OpenStudy (anonymous):

that is b!

OpenStudy (anonymous):

thanks...

OpenStudy (anonymous):

one last question and i won't bother u

OpenStudy (anonymous):

isn't is .54*.2=.108 in first case...simple p(a)p(b)

OpenStudy (anonymous):

ya...

OpenStudy (anonymous):

hey Akash, can u help me solve other problem

OpenStudy (anonymous):

1-.108...ans to part b

OpenStudy (anonymous):

0.432 for b, right?

OpenStudy (anonymous):

i need ur help..

OpenStudy (anonymous):

Akash, u there?

OpenStudy (anonymous):

part b is 1-.108 calculate this , it is not equal to .432

OpenStudy (anonymous):

Jessica Salas, president of Salas Products, is reviewing the warranty policy for her company's new model of automobile batteries. Life tests performed on a sample of 100 batteries indicated that an average life of 75 months, a standard deviation of 5 months, and a bell shaped battery life distribution. (a) between what two numbers would 95% of the batteries will fall? (b) between what two numbers would 99.7% of the batteries will fall?

OpenStudy (anonymous):

can u plz help me with this question?

OpenStudy (anonymous):

post each question as seprate

OpenStudy (anonymous):

sorry dude...

OpenStudy (anonymous):

here is the question...

OpenStudy (anonymous):

Jessica Salas, president of Salas Products, is reviewing the warranty policy for her company's new model of automobile batteries. Life tests performed on a sample of 100 batteries indicated that an average life of 75 months, a standard deviation of 5 months, and a bell shaped battery life distribution. (a) between what two numbers would 95% of the batteries will fall? (b) between what two numbers would 99.7% of the batteries will fall?